OpenStudy (anonymous):

At an interest rate of 7% compounded annually, how long will it take to double the following investments? (Round your answer to the nearest year.) (a) $200 _____ years

OpenStudy (anonymous):

JIMMY!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

hi - the real me is there today!

OpenStudy (anonymous):

i noticed

OpenStudy (anonymous):

can you help me with this? this is the formula p(1+r/n)^nt ive done problems like this but this ones missing info that i usually have given

OpenStudy (anonymous):

i'm a bit rusty at these what is n ? t = number of years right?

OpenStudy (anonymous):

r = 7/100 = 0.07

OpenStudy (anonymous):

n is 1 because it says its compounded annually and t is the years which is what i'm looking for

OpenStudy (anonymous):

this is what i get 200(1.07)^t

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

ok so the formula must be 400 = 200(1 + 0.07)^ t

OpenStudy (anonymous):

400 because you have doubled your money

OpenStudy (anonymous):

i see... i left the 400 out

OpenStudy (anonymous):

so (1 + 0.07)^ t = 400/200 = 2 1.07^t = 2

OpenStudy (anonymous):

yeah and then?

OpenStudy (anonymous):

now you can take logs: log 1.07^t = log 2 t * log 1.07 = log 2

OpenStudy (anonymous):

t = log2 / log 1.07 = 10.24

OpenStudy (anonymous):

thats 10 years

OpenStudy (anonymous):

thats right but could u put the last few steps into words please

OpenStudy (anonymous):

ok you see how i got 1.07^t = 2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then take the logs of both sides

OpenStudy (anonymous):

so we have log 2 = t * log 1.07

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

right then divide both sides by log 1.07 to get t = log 2 / log 1.07 - then use your calculator

OpenStudy (anonymous):

Ok gotcha! Thanks!

OpenStudy (anonymous):

ok :D